Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day11


Crew 102 Log Book for April 5, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: The prescriptions seem working well, crew is feeling much better today.

The rest of the Crew is doing fine.

Time Departed/Returned from EVA: 11:30 - 13:30, 14:00-15:00, 17:00-20:00

Brief Narrative of Field Mission Results:

Collected many other cyanobacteria samples. Made very good filming of the EVA for outreach documentary.

Made photos of habitat corrosion

Made last Space Florida experiment.

Mapped other 32 points on the bioerosion/corrosion study

EVA Data/Interpretations:

Cyanobacteria samples analysis ongoing

Results from Habitat corrosion visual inspection being analyzed

Engineering/Hab Maintenance:

Getting low on water, the trailer potable water tank need to be refilled

Main generator running fine after provisional fixes

Downstairs furnace still off but wall mounted heater backup is fine.

Trickle Reservoirs and recycling loop pumping running fine.

Plans for Tomorrow:

At 09:00 AM the filming crew from Arte Television will be with us again mainly for Crew interviews.

No EVA if whether is bad as forecast

Proceed on Radiotelescope data collection

Proceeding on bringing back gray water system pipeline.

Proceed on Spirulina algae growing study

Continue on ongoing psychological and physiological studies.


Support Requested:

Don Lusco for:

1) Refill water trailer tank

2) Final fixes on power generator.

3) Fix downstairs furnace

4) Bring away Telescope boxes to be sent back to the producer.

Miscellaneous:



Commander's Report

Franco Carbognani Reporting

On those days, here at the MDRS, we are at the forefront of space exploration enabling technologies studies.

Will future astronauts be able to feed themselves and breath in space habitats out of spirulina algae grown on recycled water?

The reply lies on some spirulina cultures here at the MDRS Green Hab.

Crew 102 is proud to be part of those early experiments and care a lot about those cultures which we have nicknamed "our babies". GROW SPIRULINA!, GROW!!

At 09:30 AM the filming crew from Arte Television arrived onsite. We have made a very nice EVA filming plus some indoor shots with some crew members.

Very unfortunately one of the two filming crew girls did not feel too well during the afternoon most probably because of too much desert sun. We were hoping they could share Martian Pizza during dinner with us but, because of that, they had to leave earlier. We are quite confident a nice sleep at the Hankswille Hotel will make her feel better and we hope for more nice filming tomorrow.


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Kitty provisional repair seems holding. The cooling fan is running smoothly as the whole Kitty engine.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.61

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.5

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.55

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage): 0/4, to be refilled

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 20

Trailer Potable Water Tank Level (inches from bottom): 0

Hab Potable Water Tank Level (inches from bottom): 20

Potable Water Meter Reading: 44626,6

In to GreenHab Meter Reading:


Notes/Comments:

The Trailer Potable Water Tank is now empty and ready for Don to be refilled.


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray Water System: The Trickle Reservoirs are filled with gray water

Pumps running fine and water circulating.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

Boxes containing the telescope mount and controller are ready into the Observatory in order to be handed over to Don


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):no

Tire Status: good

Odometer at end of day: 123903.8

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


ATV 3 (Viking I):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: good

Comments: this ATV does not start


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 77

Thermostat setting downstairs (Farenheit):not relevant

Actual temperature downstairs (Farenheit): 70

Heating and Ventilation Notes/Comments:

Downstairs furnace is stopped most probably for a fan not working properly.

We are using as a backup the wall mounted heater which is working fine.


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Science Report (Biology)

Lara Vimercati Reporting

During the last week the following samples of human wastewater have been filtered using X-Pack technology: Urine, greywater and 1:5 urine:greywater (this last one is a true representation of the proportions that would be produced in a space habitat). 2 replicates were prepared for each kind of sample.

Different solutions have been used to start the forward osmosis process like the X-pack provided syrup (sugar based) and different concentrations of Spirulina growth medium (Schlosser, 1982). The bags were left at 30°C for 8-12 hours for the process to be completed.

An efficiency in water yield of 36-57% was calculated. A higher efficiency is technically feasible increasing the concentration of the medium in the green port (the side of the membrane from where filtered water is to be collected) but this was not performed as the yield was sufficient for our experiment.

All the samples were inoculated with 1 ml of Spirulina culture and were incubated at 28°C in the GreenHab.

The first observations have shown that controls are growing and displaying their typical morphology, while all the cultures grown in the medium prepared with filtered water display a different color (yellow) that probably reflects a deterioration of their viability. Fully viable cultures are in fact characterized by bright green cells. This is probably due to the fact that the filtered water quality provided by X-pack is not high enough to support full growth of Spirulina under the provided conditions.

A quantitative and microscopical analysis will ultimately confirm this hypothesis.

Limitations: There is currently not suitable environment to culture cyanobacteria at MDRS. The GreenHab is not able to provide stable temperature needed to culture cyanobacteria like Spirulina and both incubators do not have a light source, also fundamental for cyanobacteria growth.


Science Report (Geology)

Kavya Manyapu Reporting

Soil Permeability, Porosity, temperature and humidity measurements

OBJECTIVE

This experiment supports the project for long life micro fuel cell power supplies for field sensors conducted by the Field and Space Robotics lab at the Massachusetts Institute of Technology. The objective of this research at MIT is to design, develop and

demonstrate the use of micro fuel cells to power sensors in extreme environments including space and terrestrial applications.  One of the terrestrial applications that this concept can be implemented is for border security where the sensors and power system will be buried in the soil (Test case is in a desert). The knowledge of soil properties such as permeability, porosity, temperature and humidity variations are required in order to implement the fuel cell powered sensor system in desert regions. Another application of this concept could be subsurface sensing for water or other parameters on planetary surfaces

BACKGROUND

Proton exchange membrane fuel cells (PEMFC) are used for missions as noted above because of they are very efficient in comparison to battery technology, and can provide power for long durations in the order of months to years because of their high energy capacity. They operate at room temperatures and are also environmentally friendly.

PEMFCs use hydrogen and air/Oxygen to produce electricity and water.

Hence, to implement this power system in the subsurface soils, the amount of Air/O2 available by the PEMs to produce enough electricity is to be learned. Soil parameters such as permeability, porosity of an area are required to completely design the fuel cell powered system.

The effects of rain water percolating into the soil might also effect the operations of the power system. Hence studying these properties will help in designing proper insulation, water and air management systems for the power systems.

MDRS, being a Mars analogue is one of the desired areas to test the soil properties. Apart from knowing the required parameters, the analogue simulation also acts as a test bed for field operations and will help develop procedures for a human space exploration of

mars/planetary surfaces involving such experiments.

Soil Permeability

Soil Permeability is the ability of water to percolate into the soil. Soil permeability gives the rate at which water/air can flow through the soil. This property is measured in unit’s cm/min (SI). This property will help in designing the water management system as well as insulating the fuel cell powered system from damage from rain water when buried in soil. Because of the high variability in soil types around the MDRS region and the unavailability of lot of equipment, just two soil types were used to study the properties of soil. Figure 1 shows the two soil types used for the experiment.


PROCEDURE

1.      A tin can of size with a diameter of 15.2 cm and length 17.3cm with both ends removed was used for this experiment.

2.      The soil was first described as best before disturbing it in any way. The following observations on the soil were performed.

a. Location (pasture, grassy field, riverbank, desert, etc.)

b. Plant material present (grass, dead leaves, etc.)

c. Soil condition (dry, moist, sandy, hard, soft, etc


3.      The can is now set on the ground and is hit hard by a hammer until the can is driven 5cm into the ground.

4.      A piece of tape is placed on the inside of the can parallel with the top edge, at 10 cm from the bottom.

5.      The distance from the bottom of the tape to the ground level is measured.

6.      The can is filled with until it reaches the level of the bottom of the tape and the time at that point is recorded.

7.      The water level will drop as the water permeates through the soil. Using a a measuring tape/ruler, the water level from the bottom of the tape is measured every 2 minutes  for soil with high permeability and

every 10 minutes for the soil with low permeability (white soil) .

This process could only be performed for a limited amount of time because of the limited availability of water in the habitat (using fresh water for the experiment was limited to 1 gallon). However, the

results obtained were sufficient.


DISCUSSION

RESULTS

Red Soil

Location        Utah Desert 38.40653ᵒ N  110.79136ᵒ W

Plant material  No Vegetation

Soil Condition  Dry, Sandy, Clay, soft

White Soil

Location        Utah Desert 38.40653ᵒ N  110.79136ᵒ W

Plant material  No Vegetation

Soil Condition  Dry, Sandy, Clay, hard

Air Temperature during experiments      Humidity Level in Air during experiments

Test 1  32.3 ᵒC 12.2%

Test 2  26.2 ᵒC 10.4%


Average Soil Permeability (cm/min)      Peak Soil permeability (cm/min)

Red Soil Type   0.86 cm/min     4.2cm/min

White Soil Type 0.026 cm/min    0.07 cm/min


The temperature and humidity experiments were done twice on two different EVAs in order to verify is the results obtained were reasonable.

OBSERVATIONS

•       The first observation is that the soil permeability rates of the white soil type are much less than the red soil. The white soil type has more clay percentage than the red soil and hence the permeability

of water in clay is much lower than sandy soil.

•       Another observation made was that while performing the experiment on the red soil type, the depth at which water percolated was measured to be 8cm. There was more of lateral movement of water than vertical movement. This might be a good result for using fuel cell power systems at depths greater than 6 inches.


Health and Safety Report

Dr. John E. Deaton Reporting

Health and Safety Officer recovering nicely now that the antibiotics have kicked in.  French film crew visited the Hab today to continue their documentary on Crew #102.  One of the film crew members appeared to have suffered heat exhaustion after a period of time outside the Hab.  She was given fluids (500 mL) and an energy bar to get her glucose levels up, and rested in one of the crew’s bedrooms.  She recovered to an extent so she could continue her interviews and will return tomorrow morning at 0900.

Other crew members have not reported anything out of the ordinary.


EVA 12 Report

Lara Vimercati Reporting

Time: 11.30-13.30

Crew Members: L.Vimercati, F.Carbognani, Y.Brodsky

Site Location: UTM 12 052 0073 4251636 (NAD 27 CONUS)

Transit Mode: ATV

Objectives: Sampling cyanobacteria that display different morphological features and live in different areas (directly on the sand forming crust-like structures and in crack inside rocks). The secondary objective was also to film crew 102 on an EVA by ARTE TV crew.

Results: ARTE TV successfully filmed all the steps involved in performing an EVA. Crew 102 members collected a large quantity of samples from different sites within a region. The samples will be stored for further analyses.

Lessons Learned: N/A.



EVA 13 Report

Dr. John E. Deaton Reporting

Time: 1330 - 1415

Crew Members: J. Deaton (EVA CDR), Y. Brodsky, A. Kapoglou

Site Location: Just outside the Hab

Transit Mode: Walking

Objectives: To conduct outreach experiment for Space Florida.  Experiment consisted of measuring the distance a radio controlled vehicle will traverse at higher altitude.  It was hypothesized by middle school team (part of outreach) that the RC car will travel further distance at altitude than at sea level.

Results: Tests completed as planned.  Crew does not have results from sea level tests, as those were conducted by the middle school team.  Crew #102 will provide altitude results so the two environments can be compared.

Lessons Learned: N/A...routine EVA


EVA 14 Report

Yuval Brodsky Reporting

Time: 1700 - 2000

Crew Members: Y. Brodsky (EVA CDR), L. Vimercati, F. Carbognani

Site Location: UTM 12 517448 4250658 (NAD27 CONUS)

Transit Mode: ATV

Objectives: Mapping of Gryphea arcuata bioerosion and corrosion along Radio ridge, with an emphasis on mapping the boundary between the two phenomena (35 predefined way-points). 

Results: All way-points were reached and both phenomena successfully mapped throughout the entire study area.

Lessons Learned: The microphones on two of three radio headsets are not working. These should be replaced, because it is frustrating, counterproductive and unsafe when comms between crewmembers on an EVA and with the HABCOM are compromised.


Journalist Report

Kavya Manyapu Reporting

Lots of things have been happening during this final week for crew 102. The French crew continues to film the events at MDRS interviewing the crew, tailing them during the EVAs, asking curious questions withregard to the dehydrated food that the crew eat, the toilet etc, etc.

Meanwhile the experiments at MDRS are slowly being finalized. Today’s first EVA commanded by F.Carbognani and accompanied by L.Vimericati and Y.Brodsky was in search of cyanobacteria. The cap-Com for this EVA was K.Manyapu. The second EVA was part of the Space Florida outreachcommanded by J.Deaton and accompanied by Y.Brodsky and A.Kapoglou and K.Manyapu served as a cap-com. The French crew were very excited by the EVAs taking place at MDRS and also had the opportunity to don the spacesuits here at MDRS to learn how the astronauts maneuver in these bulky suits. K.Manyapu conducted her soil porosity experiment at the lab and A.Kapoglou continues to work on the corrosion problem at MDRS.

The crew has been having a great time even with their busy schedule and the crew cannot believe that 2 weeks on the red planet are going to end in another 3 days.

The mission is coming to an end and the crew will return to Earth very soon. BUT, once a Martian , always a Martian!